OpenStudy (anonymous):

When you simplify each side of the equation 4(x + 3) = 5 – (x + 1), you will get the equivalent equation 4x + 12 = -x + 4. True False

OpenStudy (ajprincess):

Vat do u think?

OpenStudy (anonymous):

true

OpenStudy (anonymous):

yes if x=-8/5

OpenStudy (ajprincess):

yup it is. 4(x + 3) = 5 – (x + 1) 4x+4*3=5-x-1 4x+12=5-x-1 4x+12=-x+4

OpenStudy (anonymous):